a woman earns 15% more than her husband. Together they make 968,875 anually. Find the annual salary of the husband.
Answer by mananth(16949)

let the husband earn $x
woman earns 15% more than husband
x+1.15x=968875 ( She earns 1.15 times more than husband)
2.15x= 968875
x=968875/2.15
450,639.50 OR $450,640 eqrned by the husband